OpenStudy (anonymous):

Find the zeros of f(x)=x4-100

OpenStudy (solomonzelman):

Here, I'll do a similar example: \(\large\color{black}{ f(x)=x^4-25}\) \(\large\color{black}{ 0=x^4-25}\) \(\large\color{black}{ 0\color{red}{+25}=x^4-25 \color{red}{+25} }\) \(\large\color{black}{ 25=x^4 }\) you can first take a square root of both sides: \(\large\color{black}{ \pm 5=x^2 }\) then take a square root again: \(\large\color{black}{ - 5=x^2 ~~~\Longrightarrow~~~x=\pm i\sqrt{5}}\) \(\large\color{black}{ 5=x^2 ~~~~~~\Longrightarrow~~~x=\pm \sqrt{5}}\) so the zeros are: \(\large\color{black}{ \sqrt{5},~~-\sqrt{5},~~~~~ i\sqrt{5},~~-i\sqrt{5}}\)

OpenStudy (solomonzelman):

@americavera do you have any questions?